OpenStudy (anonymous):

Find dy/dx by implicit differentiation: 4 cos x sin y = 1

OpenStudy (mathmale):

Hello, Jay! Let's take a look at 4 cos x sin y = 1. Do you recognize that (cos x)(sin y) is a product?

OpenStudy (anonymous):

yes sir

OpenStudy (mathmale):

Jay, supposing you had (cos x)(sin x). How would you go about finding the derivative of that with respect to x? (This is for practice.) Can you comfortably apply the product rule?

OpenStudy (mathmale):

Which form of the product rule are you most comfortable with ?

OpenStudy (anonymous):

product rule is the easy to solve. but my classmate told me that he used the implicit rule sorry for my grammar i'm not fluent in english

OpenStudy (anonymous):

i think product rule is the easy way to solve

OpenStudy (mathmale):

Actually, we do have to use implicit differentiation, and are going to use it, soon. For now, please type out the form of the product rule with which you're most familiar.

OpenStudy (anonymous):

i have no idea. because i'm such a noob

OpenStudy (anonymous):

it's my assignment. pls help me :(

OpenStudy (anonymous):

-4sin(x)*sin(y) + 4cos(x)*cos(y)*dy/dx = 0

OpenStudy (mathmale):

Let me do a quick review for you: The product rule looks like this:

OpenStudy (mathmale):

\[\frac{ d }{ dx }u*v = u \frac{ dv }{ dx }+v \frac{ du }{ dx }\]

OpenStudy (mathmale):

Here we are assuming that both u and v are functions of x. Have you seen this formula before? Have you used it before?

OpenStudy (mathmale):

Jay?

OpenStudy (anonymous):

wait im thinking

OpenStudy (mathmale):

Please write that formula down, Jay. You'll need it again and again in Calculus. Label it "product rule"

OpenStudy (mathmale):

Another form of the product rule is (u*v) ' = u*v " + v*u '

OpenStudy (anonymous):

d/dx [f(x)g(x)] = f(x)g'(x) + g(x)f'(x)

OpenStudy (mathmale):

That's fine. We're on the same wavelength. Very good.

OpenStudy (mathmale):

Now, back to the example problem I gave you. y=(sin x) (cos x) is a product. We'll apply the product rule to find the derivative of this function.

OpenStudy (mathmale):

Let f(x) = sin x Let g(x) = cos x What are f '(x) and g '(x) ?

OpenStudy (anonymous):

i dont know how to solve that :(

OpenStudy (mathmale):

Is this a calculus course, or are you learning some calculus in some other course, such as Advanced Placement Math?

OpenStudy (anonymous):

sin x d/dx (cos x) + cos x d/dx (sin x) ?

OpenStudy (mathmale):

Jay, I need your complete attention if you want my help in learning this material. OpenStudy says you are "just looking around." Are you willing and able to give me your complete attention?

OpenStudy (anonymous):

yes i will give it to you

OpenStudy (anonymous):

sorry for that

OpenStudy (mathmale):

If f(x) = sin x, the derivative if f '(x) = cos x. (You really do have to know that by heart.)

OpenStudy (mathmale):

If g(x) = cox x, the derivative is g ' (x) = - sin x. I'd suggest you write all of this down for later reference, as you will see these functions, formulas and properties again and again.

OpenStudy (anonymous):

okay i get it

OpenStudy (mathmale):

Again, let's focus on finding the derivative of y= (sin x)(cos x). f(x) * g(x) You now know what f(x), g(x), f '(x) and g '(x) are. Can you apply the product rule to finding the derivative dy/dx, also using the derivatives of sin x and cos x that I've given you?

OpenStudy (anonymous):

wait give me a minute

OpenStudy (mathmale):

Write the derivative in this form:\[y ' = f * g ' + g * f '\]

OpenStudy (anonymous):

0 = (sin x ) (-sin x) + ( cos x ) ( cos x ) ?

OpenStudy (mathmale):

y ' is not zero here, but the rest of your expression is fine. Why did you think y ' = 0 ?

OpenStudy (mathmale):

Remember, this is a practice problem; it is not the actual problem that you need to solve.

OpenStudy (anonymous):

d/dx y i don't know sorry

OpenStudy (mathmale):

Actually, Jay, we are not using the power rule. We need and are using the product rule. Let me summarize this again: If y = sin x cos x then f = sin x , g = cos x, f ' = cos x and g ' = ??? (can you finish this by telling me what g ' is equal to when g = cos x?

OpenStudy (mathmale):

g ' = ?

OpenStudy (anonymous):

- sin x

OpenStudy (mathmale):

Great! Progress!

OpenStudy (mathmale):

So, Jay, if y = sin x cos x, the derivative is y ' = ( )( ) + ( ) ( ) please fill in the blanks. (This should not be new to you any longer, since we've done this already this morning.)

OpenStudy (anonymous):

(sin x) (-sin x) + (cos x) ( cos x)

OpenStudy (mathmale):

Very nice! Please label that. Your label should be y ' . Type out y ' and then copy and paste your previous result.

OpenStudy (anonymous):

y ' = (sin x) (-sin x) + (cos x) (cos x)

OpenStudy (mathmale):

Very nice. I keep asking you to type things out so that you can get the practice you need in doing so.

OpenStudy (anonymous):

thanks :)

OpenStudy (mathmale):

Now suppose you have y = 4 [ sin x cos x ]. The only difference here is that you've multiplied sin x cos x by 4. What will the derivative of y = 4 [ sin x cos x] be?

OpenStudy (mathmale):

Hint: Use the "Constant Coefficient Rule for derivatives" here.

OpenStudy (mathmale):

Don't sweat it; this is one of the easiest rules of differentiation to apply.

OpenStudy (anonymous):

y' = 0

OpenStudy (mathmale):

No, because that 4 is a constant coefficient. The derivative of 4 is zero, yes, but the derivative of 4x is definitely not zero. If y = 4 [ sinx cos x ], the derivative y ' is y ' = 4 [ ? ]

OpenStudy (mathmale):

Don't sweat this; you already have the answer (above); just copy and paste it within the brackets [ ]

OpenStudy (anonymous):

y' = 4 [ sinx * -sinx + cos x * cos x ]

OpenStudy (mathmale):

That's great. You're doing fine.

OpenStudy (mathmale):

Review: If y = sin x, (dy/dx) or y ' = cos x, right?

OpenStudy (anonymous):

yup

OpenStudy (mathmale):

Now suppose we throw out that 'x' and replace it with 'u' and note that 'u' is itself a function. The chain rule requires us to find the derivative of that 'u' with respect to the independent variable, x.

OpenStudy (mathmale):

Do you recall having heard or read about the chain rule as it applies here?

OpenStudy (anonymous):

i have no idea about in chain rule :3

OpenStudy (mathmale):

For now, you need to realize that 'u' is now a function in its own right, and that if we want to find the derivative of y = sin u, we must first differentiate with respect to u and then differentiate u itself with respect to x. Here's what this adds up to:

OpenStudy (mathmale):

Basic function y = sin x Derivative of y = sin x is y ' = cos x. Done. Finished. That's it. More complicated case: y = sin u, where 'u' is a separate function.

OpenStudy (mathmale):

Goal: find the derivative of y = sin u with respect to x. Method: First, find the derivative of sin u with respect to u: As you'll recognize from earlier, y ' = cos u.

OpenStudy (mathmale):

To finish differentiating y = sin u, you must now find the derivative of the function u and multiply it by cos u: If y = sin u , then y ' = cos u * u ' That's it. Example;

OpenStudy (mathmale):

fIND THE Derivative of y = sin (2x). nOTE THAT u = 2x here, and note that u is a separate function of x whose derivative is u ' = 2. OK with that?

OpenStudy (anonymous):

then y ' = cos 2x * 2 ?

OpenStudy (mathmale):

Perfect. Jay, you know your stuff!

OpenStudy (mathmale):

Now we're going to use implicit differentiation to find y ' if 4 cos x sin y = 1. This is the problem you posted originally.

OpenStudy (anonymous):

yeah. and i have no idea where i start

OpenStudy (mathmale):

I'll bet you do too have an idea of what to do!

OpenStudy (mathmale):

Write this out to show yourself and me what you are supposed to do: "Find y ' if 4 cos x sin y = 1 " Start by enclosing the left side and then the right side within parentheses: (4 cos x sin y ) = (1)

OpenStudy (mathmale):

Now indicate that you are to take the derivative of both sides: (4 cos x sin y ) ' = (1) '

OpenStudy (mathmale):

First, let's look at the right side. What is the derivative of a constant?

OpenStudy (mathmale):

Equivalently, what is the slope of a horizontal straight line?

OpenStudy (anonymous):

i don't how to get the slope of that line

OpenStudy (mathmale):

Review: the slope of a horizontal line is zero. The slope of the horizontal line y = 1 is zero. The slope of a line is equal to the derivative of the linear function y = 1, which again is zero. What is the derivative of y = 1? This is [ 1 ] '.

OpenStudy (mathmale):

I am saying the same thing in several different ways: The slope of a horizontal line is zero. y = 1 represents a horizontal line. Therefore, the derivative of y is zero because the derivative of 1 is zero; the derivative of a constant is zero.

OpenStudy (mathmale):

What is the right side of this equation equal to? (4 cos x sin y ) ' = (1) '

OpenStudy (anonymous):

(1) ' = 0

OpenStudy (anonymous):

still there?

OpenStudy (mathmale):

Yes

OpenStudy (mathmale):

I'm sorry OpenStudy froze up. You have the correct result for the right side of y our equation: it is zero (0).

OpenStudy (mathmale):

Therefore, (4 cos x sin y ) ' = (1) ' becomes (4 cos x sin y ) ' = 0.

OpenStudy (anonymous):

how about on the left side what should i do

OpenStudy (mathmale):

Remember, that '4' is a constant coefficient, so you can re-write (4 cos x sin y ) ' = 0 as (4) * (cos x sin y ) ' = 0. We have already found the derivative with respect to x of (cos x sin y). Could you review our work and find that derivative, and then plug it into (4)* ( ) = 0 ?

OpenStudy (mathmale):

Jay, I'm glad you're interested in communicating with me ( e. g., through Facebook), but please let's finish the solution to this problem, OK? \

OpenStudy (anonymous):

4 [ -sin x sin y + cos x cox y ] = 0 ?

OpenStudy (anonymous):

then whats next?

OpenStudy (mathmale):

Notice that I deleted my last post; I'd made a mistake there. It'll take just a minute: Let's AGAIN find the derivative of cos x sin y with respect to x: Applying both the product and chain rules, the derivative of cos x sin y with respect to x is : (cos x) (-cos y * y ') + (sin y ) (-sin x) Check that over, please, and see whether you agree or disagree or are not sure.

OpenStudy (mathmale):

I had to take the derivative of y because we are assuming that y is a separate function of x.

OpenStudy (anonymous):

but why it become (cos x ) (-cos y * y' ) + ( sin y ) (-sin x) ?

OpenStudy (mathmale):

To answer that, I'd need to know what you understand and what you don't understand. What's new about this problem, compared to my examples, is that 'y' is now a separate function of x, requiring us to use the Chain Rule.

OpenStudy (mathmale):

Point out what you do understand and what y ou do not understand about (cos x ) (-cos y * y' ) + ( sin y ) (-sin x)

OpenStudy (anonymous):

the - cos y * y'

OpenStudy (anonymous):

why it becomes negative?

OpenStudy (mathmale):

Rather than answer that, I'm asking you to show me what you think the correct result should be.

OpenStudy (anonymous):

( cos x ) (cos y * y ') + ( -sin x ) ( sin y)

OpenStudy (mathmale):

I agree with your ( cos x ) (cos y * y ') and admit I must have made a mistake there. Glad you've found it.